OpenStudy (imamermaid29):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = ____x + ____

pooja195 (pooja195):

y=mx+b Slope is given y=-4x+b Take the given coordinates and plug them in 5=4(-1)+b solve for b

OpenStudy (imamermaid29):

how would i solve for b? i suck at math

pooja195 (pooja195):

Well 4 * -1=?

OpenStudy (imamermaid29):

-4

pooja195 (pooja195):

Good 5=-4+b We want to solve for b as in get it alone add 4 to both sides 5+4=?

OpenStudy (imamermaid29):

9

pooja195 (pooja195):

Good so b=9 Any ideas on what's next?

OpenStudy (imamermaid29):

y = -4x + 9 ?

pooja195 (pooja195):

Perfect well done ^_^

OpenStudy (imamermaid29):

thank you so much!!!

pooja195 (pooja195):

:-) yw